# Ground-Compose-Reinforce: Tasking Reinforcement Learning Agents through Formal Language

### v1

@@ Abstract
Grounding language in complex perception (e.g. pixels) and action is a key challenge when building situated agents that can interact with humans via language. In past works, this is often solved via manual design of the language grounding or by curating massive datasets relating language to elements of the environment. We propose Ground-Compose-Reinforce, a neurosymbolic framework for grounding formal language from data, and eliciting behaviours by directly tasking RL agents through this language. By virtue of data-driven learning, our framework avoids the manual design of domain-specific elements like reward functions or symbol detectors. By virtue of compositional formal language semantics, our framework achieves data-efficient grounding and generalization to arbitrary language compositions. Experiments on an image-based gridworld and a MuJoCo robotics domain show that our approach reliably maps formal language instructions to behaviours with limited data while end-to-end, data-driven approaches fail.
@@ 1 Introduction
Grounded language acquisition—the way a situated agent associates language with perception and action in its environment—is a fundamental problem when building robots and other AI systems that are interfaced through language. In recent years, this problem has garnered significant interest from the AI community. One popular solution is to start with manually-designed groundings of language based on domain knowledge, such as a language-conditional reward function or success detector (e.g. [1, 2, 3, 4]). For instance, in the BabyAI benchmark [1], successful execution of instructions like “go to the red ball” can be evaluated programmatically via the environment simulator, providing a reward signal for learning language-conditioned behaviours. Such language groundings generalize to arbitrary scenarios and controlled subsets of language by design, but are hard to hand-engineer in complex, non-simulated settings based on raw perceptual inputs like pixels.
The recent advent of large language models has inspired an alternative, scalable solution to grounded language acquisition: training on diverse datasets pairing language descriptions with environment trajectories (e.g. $\pi_{0}$ [5], RT-2 [6], LIV [7], VPT [8]). While this requires no manual design of the language grounding, enormous datasets are needed to capture the intricacies of language and its corresponding range of behaviours [1, 9, 10]. Even then, fully data-driven language-grounding models are known to struggle on long-horizon or out-of-distribution instructions [11, 12, 13].
To address these issues, we propose a neurosymbolic framework called Ground-Compose-Reinforce for grounded language learning and eliciting diverse behaviours through this language. We hypothesize that compositionality plays a critical role in data-efficient and generalizable grounded language learning, by requiring only the grounding of basic concepts which can then be systematically combined to interpret complex language. To explore this hypothesis, we consider formal languages, whose syntaxes and semantics are defined by compositional rules based on an initial set of atomic symbols. Starting with no prior knowledge except for these rules, our framework: (1) grounds these atomic symbols in the environment through data; (2) composes these symbols via a formal language to express complex tasks; (3) trains an agent to solve any such task in the environment via reinforcement learning (RL) on self-generated rewards based on its grounded language understanding.
We implement this framework for Reward Machines [14], which express reward functions via temporal and logical composition over a set of binary environment properties. Given a small grounding dataset of trajectories labelled with these properties, we show how an agent can learn to ground any Reward Machine in the environment, allowing it to self-evaluate task success. To address sparse rewards, we propose a compositional reward shaping strategy that allows the agent to directly predict optimal value functions for any such task. We evaluate our agent across diverse Reward Machine tasks, including temporally-extended navigation in a 2-D gridworld and robotic manipulation in MuJoCo with raw perceptual inputs and low-level actions. We show that our approach elicits a diversity of behaviours expressible as Reward Machines, outperforming non-compositional grounding approaches according to a ground-truth interpretation of each task. Critically, we show that our approach generalizes to out-of-distribution behaviours that never appear in the grounding dataset and scales to sparse-reward settings, all while requiring no domain-specific engineering of the language grounding (e.g. of success detectors, reward functions, or symbol evaluators). More broadly, our findings highlight the important role of compositionality in grounded language learning.
@@ 2 Related Work
Recent works have proposed to learn the language grounding from diverse data, without relying on domain-specific manual design. Black et al. [5], Brohan et al. [6] and Kim et al. [16] propose vision-language-action models for real-world robotics, while Baker et al. [8] and Lifshitz et al. [13] propose similar models for playing MineCraft. Bahdanau et al. [17] and Ma et al. [7] propose approaches for learning language-conditioned reward functions directly from data. Works have also proposed directly leveraging vision-language models for rewards in vision-based environments [18, 19, 20]. These approaches do not explicitly leverage compositional language structure, and are known to struggle or hallucinate on complex, long-horizon tasks. Shi et al. [12], Yuan et al. [21], Huang et al. [22] and Ahn et al. [23] propose to use language models to decompose complex tasks at execution time. This shares motivation with our work, except we instead leverage formal language semantics to decompose tasks.
@@
Formal Languages for RL. Formal languages like Linear Temporal Logic (LTL) [24] and Reward Machines (RMs) [14] have a rich history of application towards the control [25, 26], verification [27, 28, 29], and specification of dynamical systems. More recently, they have risen in popularity in deep RL for white-box specification of reward functions [30, 31, 32, 33] and in many cases, natural language commands can be automatically translated into such formal languages [34, 35, 36, 37]. Several works show that formal languages enable compositional generalization to unseen instructions. Vaezipoor et al. [38], Kuo et al. [39] and Yalcinkaya et al. [40] train instruction-conditioned policies that zero-shot generalize to unseen instructions by training on procedurally generated LTL formulas. Qiu et al. [41], Liu et al. [42], León et al. [43] and Jackermeier and Abate [44] train transferable skills that can be invoked by a planner. Nangue Tasse et al. [45] consider how value functions and policies can be composed zero-shot for arbitrary RM tasks, but assume that all the tasks can be captured via a compact set of “goal states” that is known and enumerable. Our compositional reward shaping approach also builds on prior methods. Camacho et al. [46], Furelos-Blanco et al. [47] and Parać et al. [48] provide potential-based rewards for RMs based on the current RM state, but such methods provide no signal when the propositions themselves rarely occur. Several works propose continuous progress signals for each proposition [49, 50, 51, 52, 53, 54], but these approaches are limited to tasks with a binary success criteria. While progress signals are typically manually specified, we show how they can be learned directly from data.
@@
Nearly all formal-language-based deep RL methods are implemented with manually designed symbol evaluators (a.k.a. a labelling function) based on the environment simulator. To our knowledge, only a few works specifically avoid this assumption, but they instead depend on a manually designed reward signal. Li et al. [55, 56] consider the implications of noisy symbol groundings trained from data when using RMs. Hyde and Santos [57] and Christoffersen et al. [58] infer RMs from the reward signal as an inductive bias for RL. Umili et al. [59, 60], Kuo et al. [39], Andreas et al. [61] and Oh et al. [62] show that ungrounded formal language specifications can improve RL by providing information about the task structure. In contrast to these works, our approach does not rely on manually designed symbol evaluators or reward functions.
@@ 3.1 Reinforcement Learning
Figure 1: Four temporally extended tasks in a simple gridworld expressed as Reward Machines over the propositional vocabulary $\mathcal{AP}=\{\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\bigcirc\}$. Edges are labelled with $\langle\varphi,r\rangle$, indicating the logical condition $\varphi$ for when the corresponding transition should be followed, and the reward $r$ that is yielded as a result. Doubled circles indicate terminal states, and we omit non-rewarding self-loop edges to aid readability.
@@ 3.2 Task Specification via Reward Machines
Formal languages with sequential or temporal structure—including Reward Machines (RMs) [14], regular expressions [63], and various temporal logics [54, 38]—offer an intuitive and expressive interface for specifying tasks in RL, and support representations that capture high-level task structure and compositionality. Starting with a predefined set of atomic propositions $\mathcal{AP}$ that represent abstract, binary features of environment states, such languages can be used to express temporally extended properties. In this work, we focus on tasks specified as RMs, which subsume several formal languages of interest [46] such as LTL over finite traces [64, 65] and regular expressions.
A Reward Machine (RM) [14] is an abstract automaton that defines a reward function over the set of propositions $\mathcal{AP}$ (Figure 1). It takes as input a sequence of truth assignments $\langle\omega_{1},\omega_{2},\ldots\rangle$ over $\mathcal{AP}$, where each $\omega_{t}\in 2^{\mathcal{AP}}$ denotes the subset of $\mathcal{AP}$ that holds true at time $t$, and outputs a corresponding sequence of rewards $\langle r_{1},r_{2},\ldots\rangle$. An RM has a finite set of internal states $\mathcal{U}$ and begins in a fixed state $u_{0}$ at time $t=0$. At each step $t\geq 1$, the RM updates its state to $u_{t}\in\mathcal{U}$ and emits a reward $r_{t}$ based on the current input $\omega_{t}$ and the previous state $u_{t-1}$. This continues until the RM enters a terminal state from a designated subset $\mathcal{F}\subset\mathcal{U}$.
@@
The gridworld in Figure 1 will serve as a running example. Consider the top-left RM, which describes a task composed of three subgoals to be completed in a fixed order. The input sequence $\langle\omega_{1},\omega_{2},\ldots\rangle$ identifies the values of all propositions at each time $t$: whether the agent is at a red object, a green object, a circle, and so on. The RM state $u_{t}\in\mathcal{U}$ provides memory of which subgoals have been completed, and transitions to the next state as soon as the agent achieves the current subgoal. When a red triangle, a green object, and a blue object that isn’t a triangle are reached in that order, the RM terminates with a reward of 1.
@@
A Reward Machine (RM) is defined as a tuple $\langle\mathcal{U},u_{0},\mathcal{F},\mathcal{AP},\delta_{u},\delta_{r}\rangle$, where $\mathcal{U}$ is a finite set of states, with initial state $u_{0}\in\mathcal{U}$ and terminal states $\mathcal{F}\subset\mathcal{U}$; $\mathcal{AP}$ is a set of propositions; $\delta_{u}:\mathcal{U}\times 2^{\mathcal{AP}}\to\mathcal{U}$ is a transition function that updates the RM state based on the current truth assignment; and $\delta_{r}:\mathcal{U}\times 2^{\mathcal{AP}}\to\mathbb{R}$ is a reward function that emits a reward at each step.
As shown in Figure 1, the transition and reward functions of an RM can be intuitively and compactly specified via a set of labelled edges of the form $\langle u,u^{\prime},\varphi,r\rangle$, indicating that if a truth assignment $\omega$ satisfies $\varphi$ (denoted by $\omega\models\varphi$), then $\delta_{u}(u,\omega)=u^{\prime}$ and $\delta_{r}(u,\omega)=r$.
@@ 3.3 Grounded Interpretations
RMs specify tasks in terms of abstract symbols (e.g., $\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}}$, $\triangle$), but these symbols must be grounded in the environment to be meaningful. This is achieved via a labelling function $\mathcal{L}:\mathcal{S}\to 2^{\mathcal{AP}}$ that maps each MDP state $s$ to the set $\omega\subseteq\mathcal{AP}$ of propositions that hold true in $s$. Any such $\mathcal{L}$ grounds an RM $\mathcal{R}$ over propositions $\mathcal{AP}$, as it induces a reward sequence for any MDP trajectory $\tau$: first, states $s_{t}$ in $\tau$ are converted into truth assignments $\omega_{t}=\mathcal{L}(s_{t})$, then the RM is recursively simulated over the sequence $\langle\omega_{1},\omega_{2},\ldots\rangle$ until termination.
An RM-MDP augments a reward-free MDP with an RM and labelling function. The resulting reward function is generally non-Markovian with respect to the original state space $\mathcal{S}$, since the reward at time $t$ depends on the internal RM state $u_{t}$. One might therefore consider training a history-based policy $\pi(a_{t}|h_{t})$ to act optimally. Fortunately, this is unnecessary if the agent has oracle access to $\mathcal{L}$. In that case, the RM-MDP is Markovian over the extended state space $\mathcal{S}\times\mathcal{U}$ [14]. The agent can use $\mathcal{L}$ to compute $\omega_{t}=\mathcal{L}(s_{t})$ and recursively simulate the RM to track $u_{t}$. This allows training a policy of the form $\pi(a_{t}|s_{t},u_{t})$, where the RM state $u_{t}$ compactly encodes the history $h_{t}$ and is sufficient for optimal decision making.

### v2

# Ground-Compose-Reinforce: Grounding Language in Agentic Behaviours using Limited Data

@@ Abstract
Grounding language in perception and action is a key challenge when building situated agents that can interact with humans, or other agents, via language. In the past, addressing this challenge has required manually designing the language grounding or curating massive datasets that associate language with the environment. We propose Ground-Compose-Reinforce, an end-to-end, neurosymbolic framework for training RL agents directly from high-level task specifications—without manually designed reward functions or other domain-specific oracles, and without massive datasets. These task specifications take the form of Reward Machines, automata-based representations that capture high-level task structure and are in some cases autoformalizable from natural language. Critically, we show that Reward Machines can be grounded using limited data by exploiting compositionality. Experiments in a custom Meta-World domain with only 350 labelled pretraining trajectories show that our framework faithfully elicits complex behaviours from high-level specifications—including behaviours that never appear in pretraining—while non-compositional approaches fail.
@@ 1 Introduction
Grounding language—connecting language with perception and action within an environment—is a fundamental challenge when building robots and other agents that are interfaced through language. One popular approach to addressing this challenge is to employ a manually-designed domain-specific interpretation of language, such as a language-conditional reward function or success detector (e.g. [1, 2, 3, 4]). For instance, in the BabyAI benchmark [1], successful execution of instructions like “go to the red ball” can be evaluated programmatically in the environment simulator, providing a reward signal for learning language-conditioned behaviours. Such instances of grounded language generalize to arbitrary scenarios and controlled subsets of language by design, but are hard to hand-engineer in complex, non-simulated settings based on raw perceptual inputs like pixels.
The recent advent of large language models (LLMs) has inspired an alternative pathway to grounding language: training on diverse datasets that pair language descriptions with environment trajectories (e.g., $\pi_{0}$ [5], RT-2 [6], LIV [7], VPT [8]). While this obviates the need for manually designed language groundings, it typically requires enormous datasets in order to capture the broad scope of language usage within an environment [1, 9, 10]. For agentic applications that are data-intensive (e.g. robotics) or where access to trajectory data is limited, such data-driven language models are prone to failure on complex or out-of-distribution tasks [11, 12, 13, 14].
Figure 1: Ground-Compose-Reinforce, a lightweight framework for training RL agents directly from Reward Machine specifications, without oracles like reward functions or feature detectors. In pretraining, we learn to map propositional symbols to context-specific truth values (“is the robot holding the red block?”) and progress signals (“how close is the robot to holding the red block?”). To elicit behaviours, we prompt the agent via a Reward Machine composed of these symbols (or via natural language, if an autoformalizer is available). The agent then synthesizes its own dense reward function and interacts with the environment to learn the desired behaviour via RL.
We propose Ground-Compose-Reinforce, an end-to-end framework for training reinforcement learning (RL) agents directly from high-level task specifications, without relying on manually designed rewards or feature detectors. We represent tasks via Reward Machines (RMs) [15, 16], an automata-based task specification language that exposes temporally extended task structure in terms of a set of atomic propositional symbols. Specifically, our framework: (1) grounds these atomic symbols in the environment by pretraining on a limited dataset of labelled trajectories; (2) composes these symbols via RMs to express complex tasks; and (3) for any such task, trains an agent to solve that task via RL on self-generated dense rewards based on its learned interpretation of the RM. Overall, Ground-Compose-Reinforce enables the elicitation of a wide range of temporally and logically structured behaviours expressed as RMs (specified directly or in some cases autoformalized from natural language), requires minimal pretraining data, and does not rely on external, domain-specific oracles for training or execution. In this paper, we present the following novel contributions:
  1. 1.
    A conceptual, end-to-end framework for compositionally grounding language in behaviours based on Reward Machines and reinforcement learning (Figure 1). Critically, our framework requires minimal labelled trajectory data for pretraining, and does not require external oracles like reward functions, success detectors, or feature detectors.
  2. 2.
    A compositional reward shaping strategy for Reward Machine tasks that is critical for strong performance in Meta-World [17]. Our strategy addresses propositional sparsity, where propositions of interest (e.g. “pick up the block”) are rarely satisfied through random exploration.
  3. 3.
    Experiments across diverse Reward Machine tasks, including temporally extended gridworld navigation and robotic manipulation in Meta-World. Our approach elicits diverse behaviours in Meta-World (including out-of-distribution behaviours that never appear in the pretraining dataset) from only 350 labelled pretraining trajectories while non-compositional approaches fail.
@@ 2 Related Work
Recent works learn grounded language from data, without domain-specific manual design. Black et al. [5], Zitkovich et al. [6] and Kim et al. [19] present vision-language-action models for robotics while Baker et al. [8] and Lifshitz et al. [13] present models for playing MineCraft. Bahdanau et al. [20] and Ma et al. [7] learn language-conditioned reward functions from data. Works have also directly leveraged vision-language models for rewards in vision-based environments [21, 22, 23], but typically do not leverage language compositionality and are prone to failure on complex or out-of-distribution tasks. Shi et al. [12], Yuan et al. [24], Huang et al. [25] and Ahn et al. [26] decompose complex tasks at execution time via language models. While this shares motivation with our work, we represent tasks via RMs and exploit compositional task structure for both training and execution.
@@
Formal Languages for RL. Formal languages like Linear Temporal Logic (LTL) [27] and other associated formal structures [^1] Henceforth “formal languages,” for ease of exposition. such as RMs [16] have a rich history of application in the control [28, 29], verification [30, 31, 32], monitoring, and synthesis of dynamical systems. Recently, they have risen in popularity in deep RL for white-box specification of rich temporally extended reward functions [15, 33, 34, 35] and in a number of cases can be automatically generated from natural language commands (e.g., [36, 37, 38, 39]). Several works show that formal languages enable compositional generalization to unseen instructions. Vaezipoor et al. [40], Kuo et al. [41] and Yalcinkaya et al. [42] train instruction-conditioned policies that zero-shot generalize to unseen instructions by training on procedurally generated LTL formulas. Qiu et al. [43], Liu et al. [44], León et al. [45] and Jackermeier and Abate [46] train transferable skills that can be invoked by a planner. Nangue Tasse et al. [47] consider how value functions and policies can be composed zero-shot for arbitrary RM tasks, but assume that all the tasks can be captured via a finite, predetermined set of goal states. Our compositional reward shaping approach also builds on prior methods. Camacho et al. [34], Furelos-Blanco et al. [48] and Parać et al. [49] provide potential-based rewards for RMs based on the current RM state, but such methods provide no signal when target propositions rarely occur. Several works propose continuous progress signals for each proposition [33, 50, 51, 52, 53, 54], but these approaches are limited to tasks with a binary success criterion. While progress signals are typically manually specified, we show how they can be learned directly from data.
@@
Nearly all formal-language-based deep RL methods assume access to an external evaluator of symbolic features (a.k.a. a labelling function). To our knowledge, only a few works specifically avoid this assumption, but they instead depend on an external reward signal. Li et al. [55, 56] consider the implications of noisily grounding symbols when using RMs. Hyde and Santos [57] and Christoffersen et al. [58] infer RMs from the reward signal as an inductive bias for RL. Umili et al. [59, 60], Kuo et al. [41], Andreas et al. [61] and Oh et al. [62] show that ungrounded formal specifications can improve RL by providing information about the task structure. In contrast to these works, our approach does not rely on an external symbol evaluator or reward function.
@@ 3.1 Reinforcement Learning
Figure 2: Four temporally extended tasks in a gridworld expressed as Reward Machines over the propositions $\mathcal{AP}=\{\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\$bigcirc\}. An edge labelled $\langle\varphi,r\rangle$ indicates the logical condition $\varphi$ for when the corresponding transition should be followed, and the reward $r$ that is yielded as a result. Doubled circles indicate terminal states, and we omit non-rewarding self-loop edges to aid readability.
@@ 3.2 Task Specification via Reward Machines
Formal languages with sequential or temporal structure—including RMs [16], regular expressions [63], and various temporal logics [54, 40]—offer an intuitive and expressive interface for specifying tasks in RL while supporting representations that capture compositional task structure. Starting with a predefined set of atomic propositions $\mathcal{AP}$ that represent abstract, binary features of environment states, such languages can be used to express temporally extended properties or reward functions. In this work, we focus on tasks specified as RMs, which subsume several formal languages of interest [34] such as LTL over finite traces [64, 65] and regular expressions.
An RM is an automaton that captures the structure of a reward function over the abstract vocabulary $\mathcal{AP}$ (Figure 2). It takes as input a sequence of truth assignments $\langle\omega_{1},\omega_{2},\ldots\rangle$ over $\mathcal{AP}$, where each $\omega_{t}\in 2^{\mathcal{AP}}$ denotes the subset of $\mathcal{AP}$ that holds true at time $t$, and outputs a corresponding sequence of rewards $\langle r_{1},r_{2},\ldots\rangle$. An RM has a finite set of internal states $\mathcal{U}$ and begins in a fixed state $u_{0}$ at time $t=0$. At each step $t\geq 1$, the RM updates its state to $u_{t}\in\mathcal{U}$ and emits a reward $r_{t}$ based on the current input $\omega_{t}$ and the previous state $u_{t-1}$. This continues until the RM enters a terminal state from a designated subset $\mathcal{F}\subset\mathcal{U}$.
@@
The gridworld in Figure 2 will serve as a running example. Consider the top-left RM, which describes a task composed of three subgoals to be completed in a fixed order. When given an input sequence $\langle\omega_{1},\omega_{2},\ldots\rangle$ that identifies the values of all propositions at each time $t$ (whether the agent is at a red object, a green object, a circle, and so on), the RM state $u_{t}\in\mathcal{U}$ tracks which subgoals have been completed and transitions to the next RM state as soon as the agent achieves the current subgoal. When a red triangle, a green object, and a blue object that isn’t a triangle are reached in that order, the RM terminates with a reward of 1.
@@
A Reward Machine $\mathcal{R}$ is defined as a tuple $\langle\mathcal{U},u_{0},\mathcal{F},\mathcal{AP},\delta_{u},\delta_{r}\rangle$, where $\mathcal{U}$ is a finite set of states, with initial state $u_{0}\in\mathcal{U}$ and terminal states $\mathcal{F}\subset\mathcal{U}$; $\mathcal{AP}$ is a set of propositions; $\delta_{u}:\mathcal{U}\times 2^{\mathcal{AP}}\to\mathcal{U}$ is a transition function that updates the RM state based on the current truth assignment; and $\delta_{r}:\mathcal{U}\times 2^{\mathcal{AP}}\to\mathbb{R}$ is a reward function that emits a reward at each step.
As shown in Figure 2, the transition and reward functions of an RM can be intuitively and compactly specified via a set of labelled edges of the form $\langle u,u^{\prime},\varphi,r\rangle$, indicating that if a truth assignment $\omega$ satisfies the formula $\varphi$ (denoted by $\omega\models\varphi$), then $\delta_{u}(u,\omega)=u^{\prime}$ and $\delta_{r}(u,\omega)=r$.
@@ 3.3 Grounded Interpretations
RMs express tasks in terms of abstract symbols (e.g., $\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}}$, $\triangle$), but these symbols must be grounded in the environment to be meaningful. This is achieved via a labelling function $\mathcal{L}:\mathcal{S}\to 2^{\mathcal{AP}}$ that maps each MDP state $s$ to the set $\omega\subseteq\mathcal{AP}$ of propositions that hold true in $s$. For an RM $\mathcal{R}$ over propositions $\mathcal{AP}$, any such $\mathcal{L}$ also grounds $\mathcal{R}$ in the environment by inducing a reward sequence for any MDP trajectory $\tau$: first, states $s_{t}$ in $\tau$ are converted into truth assignments $\omega_{t}=\mathcal{L}(s_{t})$, then the RM is simulated over the sequence $\langle\omega_{1},\omega_{2},\ldots\rangle$ to generate a sequence of rewards until termination.
An RM-MDP augments an environment (represented by a reward-free MDP) with a concrete reward function captured by an RM and labelling function. The resulting reward function is generally non-Markovian with respect to MDP states $\mathcal{S}$, since the reward at time $t$ depends on the internal RM state $u_{t}$. While one might consider expressing optimal behaviours via a history-based policy $\pi(a_{t}|h_{t})$, this is unnecessary if the agent has oracle access to $\mathcal{L}$ since RM-MDPs are Markovian over the extended state space $\mathcal{S}\times\mathcal{U}$ [16]. The agent can use $\mathcal{L}$ to compute $\omega_{t}=\mathcal{L}(s_{t})$ and recursively simulate the RM to track $u_{t}$. Thus, it is typical to express policies in the form $\pi(a_{t}|s_{t},u_{t})$, where the RM state $u_{t}$ compactly encodes the history $h_{t}$ and is sufficient for optimal decision making.
@@
Running Example.
The RM at the top left of Figure 2 captures the high-level structure of a multi-stage task. To map environment trajectories into concrete rewards for this RM, we need a labelling function $\mathcal{L}:\mathcal{S}\to 2^{\mathcal{AP}}$ that connects abstract propositions like $\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}}$ and $\triangle$ to environment states.